OpenStudy (anonymous):

H(x)=(x^2)+1 and K(x)=-(x^2)+4. If K(H)=0, what are the roots/solutions? a. ±i√3 b. ±i c. ±2 d. ±1 e. ±1, ±i√3

OpenStudy (jdoe0001):

\(\bf H(x)=x^2+1\qquad K(x)=-x^2+4 \\ \quad \\ K(H)=-(x^2+1)^2+4\qquad thus \quad K(H)=0\to 0=-(x^2+1)^2+4\) solve for "x"